OpenStudy (anonymous):

Out of 40 students, 14 are taking English and 29 are taking Chem. If 5 students are in both, how many students are in neither class

OpenStudy (anonymous):

3

OpenStudy (anonymous):

can you show your work so i can understand it? thanks

OpenStudy (anonymous):

@Zekarias

OpenStudy (anonymous):

So we know, by the principle of Inclusion exclusion: \[ |A\cup B|=|A|+|B|-|A\cap B| \]So, we wish to find the remaining of the fourty such students, we say: \[ |A\cup B|=14+29-5=38 \]So, the total number of students is 40, and the students taking at least one of Chem OR English is 38. So we must have 2 students not in a class.

OpenStudy (anonymous):

First find the number students who are taking both subject : 14+29-5 = 37 Then subtract it from the total : 40-37 = 3

OpenStudy (anonymous):

Lolwolf what is the value for 14+29-5

OpenStudy (anonymous):

You tell me. 29-5=24 And 24+14=38.

OpenStudy (anonymous):

right, thanks

OpenStudy (anonymous):

thanks guys!!!